OpenStudy (anonymous):

The attendance at 3 football games fluctuated above and below the average according to the following table. What was the net change from the average attendance after these three games? Game 1: -583 Game 2: 503 Game 3: -93

OpenStudy (anonymous):

how do i solve??

OpenStudy (anonymous):

add the three up and divide by 3. So ((-583)+503+(-93))/3=-173/3=57.666666

OpenStudy (anonymous):

-57.66666666666

OpenStudy (anonymous):

oh ok i wasn't sure how to make an equation but i see how

OpenStudy (anonymous):

it would be a negative answer right?